Suppose that quiz scores in a beginning statistics class have a mean of 7.6 with a standard deviation of 0.50.5. Using Chebyshev's Theorem, what is the minimum percentage of quiz scores between 6.1 and 9.1? Round your answer to one decimal place.
Answer #1

6.1 = 7.6 - 3*0.5 that is 3 standard deviations less than the mean.
9.1 = 7.6 + 3*0.5 that is 3 standard deviations more than the mean.

According to chebyshev's theorem, at least 1 - 1/k2 of the observations lies within k standard deviations of the mean.

Therefore at least 1 - 1/32 = 8/9 that is 88.9% of the observations lies within 3 standard deviations of the mean.